# Unified Approach to Power-Efficiency Trade-Off of Generic Thermal Machines

### v1

@@ Abstract
Acknowledgment.—This work is supported by the National Natural Science Foundation of China for support under grant No. 12305037 and the Fundamental Research Funds for the Central Universities under grant No. 2023NTST017.
Appendix A Thermodynamic geometry approach to finite-time Carnot cycle with ideal gas
@@
In the following, we specifically demonstrate how to use thermodynamic geometric approach to determine the dissipation ratio $\delta$ in a finite-time Carnot cycle with an ideal gas as the working substance. When the gas with a molar amount of $n$ and volume $V$ is driven slowly in a hot reservoir of temperature $T_{\rm{h}}$ (finite-time high-temperature isothermal process of duration $\tau_{\rm{h}}$), the gas pressure $p$ reads [12]
@@
$\mathcal{L}_{\rm{h}}$ $=\int_{0}^{\tau_{1}}\sqrt{\dot{W}_{\rm{ex}}(\textit{t})}\rm{d}\textit{t}=\int_{0}^{\tau_{1}}\sqrt{\textit{p}_{\rm{ex}}(\textit{t})\dot{\textit{V}}}\rm{d}\textit{t}$
@@
where $p_{\rm{ex}}\equiv|p-p_{\rm{q}}|$ has been used. Thus, the corresponding dissipation in this process follows as [50]
@@
Since $\kappa_{\rm{h}}/\kappa_{\rm{c}}\in(0,\infty)$, one has $\delta\in(0,1)$. Especially, in the symmetric case with $\kappa_{\rm{h}}/\kappa_{\rm{c}}=1$, it follows from Eq. (13) that $\eta_{\rm{MP}}=1-\sqrt{1-\eta_{\rm{C}}}$, which is exactly the Curzon-Ahlborn efficiency [15]
Appendix B Upper and lower bound for efficiency at given power
The range of efficiency at maximum power indicates a varying trade-off relation in heat engines as a function of the dissipation ratio, allowing us to establish tighter overall upper and lower bounds of efficiency at given power from Eq. (12)
@@ Appendix B Upper and lower bound for efficiency at given power
Under the condition that the power and thermodynamic parameters are fixed, $\tilde{\eta}_{\pm}$ delineate the upper and lower bounds of the low-dissipation region. Generally, precise knowledge of the heat engine parameters is elusive; hence, deriving a universal upper/lower bounds that are valid for all $\delta$ within this range is of practical significance. Next, we prove the monotonic increasing nature of the tight upper and lower bounds for efficiency, as a function of a given power with a variable parameter $\delta$.
@@
where $a_{-}=\tilde{P}+2\sqrt{1-\tilde{P}}-2,~{}b_{-}=4-4\sqrt{1-\tilde{P}},~{}x=\delta\eta_{\rm{rev}}$. The sign of the above equation is determined by $a_{-}x^{2}+b_{-}x-4$. Notably, it is straightforward to observe that $a_{-}\leq 0$, with equality holding only when $\tilde{P}=1$. Furthermore, under the condition $\Delta=b_{-}^{2}+16a_{-}=0$, we find that $a_{-}x^{2}+b_{-}x-4\leq 0$ for $\tilde{P}\neq 1$. When $\tilde{P}=1$ both $a_{-}$ and $b_{-}$ vanish, reducing the quadratic expression as $-4$. We conclude that $\frac{\rm{d}\tilde{\eta}_{-}}{\rm{d}\delta}>0$ for $\delta\in[0,1]$. Thus, the overall lower bound is reached at $\delta=0$, substituting which into Eq. (29) yields $\tilde{\eta}_{\rm{L}}=(1-\sqrt{1-\tilde{P}})/2$. Similarly, for

### v2

@@ Acknowledgment
This work is supported by the National Natural Science Foundation of China under grant No. 12305037, and the Fundamental Research Funds for the Central Universities under grant No. 2023NTST017.
Appendix A Thermodynamic geometry of ideal gas
@@
In the following, we specifically demonstrate how to use the thermodynamic geometric approach to determine the dissipation ratio $\delta$ in a finite-time Carnot cycle with an ideal gas as the working substance. When the gas with a molar amount of $n$ and volume $V$ is driven slowly in a hot reservoir of temperature $T_{\rm{h}}$ (finite-time high-temperature isothermal process of duration $\tau_{\rm{h}}$), the gas pressure $p$ reads [15]
@@
$\mathcal{L}_{\rm{h}}$ $=\int_{0}^{\tau_{1}}\sqrt{\dot{W}_{\rm{ex}}(\textit{t})}\rm{d}\textit{t}=\int_{0}^{\tau_{1}}\sqrt{-\textit{p}_{\rm{ex}}(\textit{t})\dot{\textit{V}}}\rm{d}\textit{t}$
@@
where $p_{\rm{ex}}\equiv p-p_{\rm{q}}$ has been used. Thus, the corresponding dissipation in this process follows as [47]
@@
Since $\kappa_{\rm{h}}/\kappa_{\rm{c}}\in(0,\infty)$, one has $\delta\in(0,1)$. Especially, in the symmetric case with $\kappa_{\rm{h}}/\kappa_{\rm{c}}=1$, it follows from the EMP equation of $\alpha=1$ that $\eta_{\rm{MP}}=1-\sqrt{1-\eta_{\rm{C}}}$, which is exactly the Curzon-Ahlborn efficiency [22].
Appendix B Upper bound of efficiency from Eq. (15)
The range of EMP indicates a varying trade-off relation in heat engines as a function of the dissipation ratio, allowing us to establish tighter overall upper and lower bounds of efficiency at given power from Eq. (15)
@@
Under the condition that the power and thermodynamic parameters are fixed, $\tilde{\eta}_{\pm}$ delineate the upper and lower bounds of the low-dissipation region. Generally, precise knowledge of the heat engine parameters is elusive; hence, deriving universal upper/lower bounds that are valid for all $\delta$ within this range is of practical significance. Next, we prove the monotonic increasing nature of the tight upper and lower bounds for efficiency, as a function of a given power with a variable parameter $\delta$.
@@
where $a_{-}=\tilde{P}+2\sqrt{1-\tilde{P}}-2,~b_{-}=4-4\sqrt{1-\tilde{P}},~x=\delta\eta_{\rm{rev}}$. The sign of the above equation is determined by $a_{-}x^{2}+b_{-}x-4$. Notably, it is straightforward to observe that $a_{-}\leq 0$, with equality holding only when $\tilde{P}=1$. Furthermore, under the condition $\Delta=b_{-}^{2}+16a_{-}=0$, we find that $a_{-}x^{2}+b_{-}x-4\leq 0$ for $\tilde{P}\neq 1$. When $\tilde{P}=1$ both $a_{-}$ and $b_{-}$ vanish, reducing the quadratic expression as 4-4. We conclude that $\frac{\rm{d}\tilde{\eta}_{-}}{\rm{d}\delta}>0$ for $\delta\in[0,1]$. Thus, the overall lower bound is reached at $\delta=0$, substituting which into Eq. (34) yields $\tilde{\eta}_{\rm{L}}=(1-\sqrt{1-\tilde{P}})/2$. Similarly, for